Two cars, A and B, are traveling with the same speed of 40.0 m/s, each having started from rest. Car A has a mass of 1200 kg, and car B has a mass of 2000 kg. Compared to the work required to bring car A up to speed, how much additional work is required to bring car B up to speed?
1
Expert's answer
2018-07-09T15:33:08-0400
The additional work is
W=K_B-K_A
W=m_B v^2/2-m_A v^2/2=(m_B-m_A)v^2/2
W=(2000-1200) 〖40〗^2/2=640000 J=640 kJ.